OpenStudy (anonymous):

What are the possible number of positive real, negative real, and complex zeros of f(x) = -7x4 - 12x3 + 9x2 - 17x + 3?

ganeshie8 (ganeshie8):

Descartes rule of signs

OpenStudy (jdoe0001):

hmmm, let's take a peek at your equation

OpenStudy (anonymous):

what descartes rule of signs?

OpenStudy (jdoe0001):

as ganeshie8 already pointed out, Descartes Rule of signs meaning how many times does jumping from one term to the next, changes signs let's see -7x4 - 12x3 + 9x2 - 17x + 3 - - + - +

OpenStudy (jdoe0001):

look at the signs of the coefficients, how many times did they CHANGE, that is from negative to positive or positive to negative if you have negative to negative, there's no CHANGE

OpenStudy (anonymous):

so they changed 3 times ?

OpenStudy (jdoe0001):

-7x4 - 12x3 + 9x2 - 17x + 3 - - + - + ^ ^ ^ ^ no yes yes yes

OpenStudy (jdoe0001):

yes, they CHANGED 3 times, that means 3 OR 3-2 = 1 or 3 or 1 positive real roots

OpenStudy (anonymous):

Positive Real: 3 or 1 Negative Real: 1 Complex: 2 or 0

OpenStudy (anonymous):

so this is my answer?

OpenStudy (jdoe0001):

now let's check how many negative ones, by using f(-x)

OpenStudy (jdoe0001):

-7x^4 + 12x3 + 9x2 + 17x + 3 - + + + +

OpenStudy (jdoe0001):

so it CHANGES only between the 1st term and the 2nd one, ONCE so that means, 1 negative real root

OpenStudy (jdoe0001):

so, the degree of the polynomial is 4, since that's the leading exponential \("7x^4"\) that means that there'll be 4 roots 3 positive and 1 negative, 3+1 = 4, so 0 complex if 1 positive and 1 negative, 1+1 = 2, so 2 complex

OpenStudy (jdoe0001):

so, your answer is right

OpenStudy (anonymous):

okay! got it! thank you!

OpenStudy (jdoe0001):

yw